OpenStudy (anonymous):

Long division for x^4 - 2x^3 + 3x^2 - 4x + 6 by x^2 + 2x -1?

OpenStudy (anonymous):

\[x ^{2} + 2x - 1 \div x ^{4} - 2x ^{3} + 3x ^{2} - 4x + 6\]

OpenStudy (triciaal):

switch

OpenStudy (anonymous):

\[x ^{4} - 2x^3 + 3x^2 - 4x + 6 \div x^2 + 2x - 1\]

OpenStudy (triciaal):

yes

OpenStudy (calculusxy):

\[\ \large \frac{ x^4 - 2x^3 + 3x^2 - 4x + 6 }{ x^2 +2x - 1 }\] when u r dividing, you are basically subtracting the exponents with the same root for example \[\large \frac{ x^6 }{ x^2 } = x^4\]

OpenStudy (calculusxy):

and if we have \[\large 2x \] then the exponent is 1 \[\large 2x^1\]

OpenStudy (anonymous):

Would x^4 end up divided (subtracting, I guess) by x^2, turning into x^2?

OpenStudy (calculusxy):

yes

OpenStudy (anonymous):